OpenStudy (anonymous):

hi guys so i am studiing 7th grade amatha nd i was wondeering what a linear function is? thanks answer fast please

OpenStudy (anonymous):

\[y=mx+b\] this is linear function

OpenStudy (anonymous):

its any equation that follows the format that imran just posted and when graphed it is a straight line.

OpenStudy (anonymous):

what do the leeters mean?

OpenStudy (anonymous):

i mean the letters srry:)

OpenStudy (anonymous):

y=output, m is the slope of the line, and x is the input

OpenStudy (anonymous):

keep explaining thankls

OpenStudy (anonymous):

(this is based off of a tchart. like so... y x 0 1 1 2 2 3 and so on.

OpenStudy (anonymous):

aww stupid open study... ill draw one for you.

OpenStudy (anonymous):

here is a t-chart. would you like me to graph it so you can see what it would look like?

OpenStudy (anonymous):

yes thankyou so much you are so awesome

OpenStudy (anonymous):

haha no problem. the equation would be y= (1/1)x +1 this is because the difference between each point is rise one, run one. this is slope. +1 is where the line intersects the y axis.

OpenStudy (anonymous):

but you can simplify this to just be y=x+1

OpenStudy (anonymous):

this makes sense because if you look at the chart, you can see that every x value is 1 greater than the y value

OpenStudy (anonymous):

yes i kinda understand i just dont get the y=x+ 1 part

OpenStudy (anonymous):

well you'll understand more as you learn more of algebra. dont worry about it right now. its just the equation fro finding y given a pattern and the x value.